# BAD-NEUS: Rapidly converging trajectory stratification

### v2

@@ VII Conclusions
Trajectory stratification methods enable enhanced sampling for estimating both thermodynamic (equilibrium) and kinetic (nonequilibrium) statistics. Here, we introduced a new trajectory stratification method, BAD-NEUS, which converges faster than existing ones. We show that our method is a natural generalization of NEUS and EM, which in turn can be formulated as extensions of WE. In the process, we show that WE as originally formulated [1] converges no faster than unbiased dynamics. The key modification that we introduce is the insertion of an approximation to the steady-state distribution before the resampling step. Importantly, we design this approximation algorithm so that it preserves the steady-state distribution of the dynamics, and therefore does not introduce systematic errors into the algorithm in the large-data limit.
@@ Acknowledgements.
This work was supported by National Institutes of Health award R35 GM136381 and National Science Foundation award DMS-2054306. This work was completed in part with resources provided by the University of Chicago Research Computing Center. “Beagle-3: A Shared GPU Cluster for Biomolecular Sciences” is supported by the National Institutes of Health (NIH) under the High-End Instrumentation (HEI) grant program award 1S10OD028655-0.
@@ Appendix A Derivation of (16)
Here we derive (16) by modifying the proof from Ref. 56. Let $Z_{t}$ be a Markov process, $\mathcal{T}_{t}$ be operator defined in (41), and $S_{n}=\sum_{t=0}^{n-1}g(Z_{t})$ be the partial sum of a function $g$. Let $\omega(z)$ satisfy the linear problem
@@
We refer the reader to Ref. 56 for a more technical discussion and conditions under which the relevant expectations are bounded such that we may apply the optional stopping theorem. Then, for any stopping time $T$ with $\operatorname{\mathbbm{E}}[T]<\infty$,

### v3

@@ VII Conclusions
Trajectory stratification methods enable enhanced sampling for estimating both thermodynamic (equilibrium) and kinetic (nonequilibrium) statistics. Here, we introduced a new trajectory stratification method, BAD-NEUS, which converges faster than existing ones. We show that our method is a natural generalization of NEUS and EM, which in turn can be formulated as extensions of WE. In the process, we show that WE as originally formulated [1] converges no faster than unbiased dynamics. The key modification introduced in NEUS and elaborated here is the insertion of an approximation to the steady-state distribution before the resampling step. Importantly, we design this approximation algorithm so that it preserves the steady-state distribution of the dynamics, and therefore does not introduce systematic errors into the algorithm in the large-data limit.
@@ Acknowledgements.
We thank Daniel Zuckerman for discussions concerning the manuscript. This work was supported by National Institutes of Health award R35 GM136381 and National Science Foundation award DMS-2054306. This work was completed in part with resources provided by the University of Chicago Research Computing Center. “Beagle-3: A Shared GPU Cluster for Biomolecular Sciences” is supported by the National Institutes of Health (NIH) under the High-End Instrumentation (HEI) grant program award 1S10OD028655-0.
@@ Appendix A Derivation of (16)
Here we derive (16) by modifying the proof from Ref. 55. Let $Z_{t}$ be a Markov process, $\mathcal{T}_{t}$ be operator defined in (41), and $S_{n}=\sum_{t=0}^{n-1}g(Z_{t})$ be the partial sum of a function $g$. Let $\omega(z)$ satisfy the linear problem
@@
We refer the reader to Ref. 55 for a more technical discussion and conditions under which the relevant expectations are bounded such that we may apply the optional stopping theorem. Then, for any stopping time $T$ with $\operatorname{\mathbbm{E}}[T]<\infty$,